# Reconstructing Galaxy Cluster Mass Maps using Score-based Generative Modeling

### v1

@@ 4.3 Cross Correlations
Figure 9: Normalized cross-correlation coefficients between the simulation and generated outputs (blue) of gas (left) and dark matter (right), for both high mass clusters (top) and low mass clusters (bottom). We expect near-perfect agreement of unity on large scales (small $k$), and a monotonically decreasing behavior as we move to smaller scales (large $k$). By comparing with the cross-coefficients between the simulated outputs and SZ (orange) and X-ray (purple), we see that our model is able to learn a much stronger nonlinear correlation than that of a scale-dependent linear bias model.
@@
high mass Gas $1.004^{+0.003}_{-0.005}$ $1.008^{+0.005}_{-0.007}$ $0.949^{+0.028}_{-0.034}$ $0.998^{+0.002}_{-0.002}$
Dark Matter $0.888^{+0.115}_{-0.098}$ $1.008^{+0.008}_{-0.011}$ $0.821^{+0.149}_{-0.131}$ $0.994^{+0.006}_{-0.005}$
low mass Gas $1.023^{+0.025}_{-0.031}$ $1.004^{+0.005}_{-0.007}$ $0.932^{+0.050}_{-0.052}$ $0.998^{+0.002}_{-0.002}$
Dark Matter $0.928^{+0.092}_{-0.081}$ $1.030^{+0.005}_{-0.006}$ $0.832^{+0.143}_{-0.123}$ $0.995^{+0.005}_{-0.005}$
@@
Figure 9 shows the pairwise cross correlation between the simulated output (ground truth) with the inputs and the generated output in blue. This curve gives us a metric on how well the reconstructed mass maps, across all scales, agree with the simulation mass maps. Plotted in the orange and purple curves are the correlation between the simulated SZ and X-ray inputs with the simulated outputs, respectively, which is a proxy for the best performance a linear bias model can do to map the inputs to the output mass maps. This effectively serves as a baseline model for mapping the observables to the underlying densities.
For the correlated simulation inputs and outputs, the model behaves similarly for both high-mass and low-mass clusters: the gas is better correlated with the input SZ and X-ray observables than the dark matter since the construction of the gas maps is much more closely aligned with the actual observables than the underlying dark matter maps, which have a much more complex relationship. More specifically, for the gas maps, SZ correlations are stronger than those of X-ray, which agree with Equations 6 and 7: because SZ is proportional to pressure and density while X-ray is proportional to the square of density, we can expect the X-ray signal to be less correlated as the square will be much more sensitive to the small-scale clumpiness of the gas maps. Finally, the diffusion results in blue show that our model is able to learn a much stronger correlation across all scales. de Andres et al. (2024) reconstructs the generated and simulated power spectrum and shows agreement by comparing the relative error, but does not compute $b(k)$ or $r(k)$, which we focus on to show the improvements of our model over linear models in the spectral domain.
In Table 1, we summarize the average values of Figures 8 and 9. Our bias values are in agreement with unity across all scales for both gas and dark matter maps except for high-mass gas maps at the large scale, which is within $0.001$ agreement with unity. Our cross-correlation values are close to unity at large scales, and around between $0.8$ to $0.95$ for small scales, indicating that our model is able to reproduce the structure of clusters very well at these two different scale lengths.
@@ 5 Conclusion
Our diffusion model is able to generate highly realistic and accurate realizations of gas and dark matter maps. The samples from the model converge since a 50-sample mean is able to achieve a $<5\%$ 2-sample variance for our test dataset. In addition, because the dataset is normalized by a global maximum and minimum so as to preserve relative intensities between clusters, the model is able to accurately predict the matter distribution across differently-sized clusters. Another indication that our model is able to distinguish between differently-sized clusters is that it is able to reconstruct the mean and spread of both the gas and dark matter density profiles to within 5% on all spatial scales and all cluster sizes. Finally, in the spectral domain, the model achieves a close-to-unity agreement for the bias coefficients and high cross-correlation coefficients between the simulated and generated density maps, implying that our model learns a strong nonlinear mapping that can accurately probe both large and small scale cluster structures.
@@
Future work can be done to improve the reconstruction of the mass maps in both the spatial and spectral domains. Currently, our loss is only the score-matching term, which does not include any information on the physical constraints of our data. We can utilize physics-informed losses during learning, where we augment our loss with extra terms, such as constraints on the total mass of the cluster or regularization terms on the bias and cross-correlation coefficients at different scale lengths. Improvements can also be made on the data end: by adding structured noise to our dataset, we can increase the robustness of our model towards imperfect observations.
@@
The success of our diffusion model to accurately predict gas and dark matter realizations by sampling from a learned distribution has huge implications in real-data inference: our model can be eventually fine-tuned and generalized to not only take in real SZ and X-ray observables, but also additional channels such as weak lensing measurements that can provide more insight towards the internal cluster dynamics. These diffusion models can ultimately be used to predict unknown gas and dark matter distributions of real galaxy clusters of interest.
@@
We thank Jun-Yan Zhu for discussions on generative modeling, and Nick Gnedin, Yuuki Omori, and Qirong Zhu for discussions on HYPER. We also thank Alexandre Adam for making his implementation of score models publicly available. Finally, we thank Ashley Villar for guidance on parts of this project during the Spring 2024 ASTRON205 course at Harvard University. The HYPER simulation was run on Endeavour at the NASA Advanced Supercomputing (NAS) Division. The machine learning work was run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University. This project is supported by NASA grant 80NSSC22K0821.

### v2

@@ 4.3 Cross Correlations
Figure 9: Normalized cross-correlation coefficients between the simulation and generated outputs (blue) of gas (left) and dark matter (right). Coefficients for high mass clusters ($6\times 10^{14}\leq M_{200m}/M_{\odot}\leq 1\times 10^{15}$) are shown on the top, and those of low mass clusters ($2\times 10^{14}\leq M_{200m}/M_{\odot}<6\times 10^{14}$) are shown on the bottom. We expect near-perfect agreement of unity on large scales (small $k$), and a monotonically decreasing behavior as we move to smaller scales (large $k$). By comparing with the cross-coefficients between the simulated outputs and SZ (orange) and X-ray (purple), we see that our model is able to learn a much stronger nonlinear correlation than that of a scale-dependent linear bias model.
@@
Figure 9 shows the pairwise cross correlation between the simulated output (ground truth) with the inputs and the generated output in blue. This curve gives us a metric on how well the reconstructed mass maps, across all scales, agree with the simulation mass maps. Plotted in the orange and purple curves are the correlation between the simulated SZ and X-ray inputs with the simulated outputs, respectively, which is a proxy for the performance of a linear bias model, which is effectively a baseline model for mapping the observables to the underlying densities. This graph serves to compare the spectral reconstruction performance of our model with the optimal performance of a linear bias model.
For the correlated simulation inputs and outputs, the model behaves similarly for both high-mass and low-mass clusters: the gas is better correlated with the input SZ and X-ray observables than the dark matter since the construction of the gas maps is much more closely aligned with the actual observables than the underlying dark matter maps, which have a much more complex relationship. More specifically, for the gas maps, SZ correlations are stronger than those of X-ray, which agree with Equations 6 and 7: because SZ is proportional to pressure and density while X-ray is proportional to the square of density, we can expect the X-ray signal to be less correlated as the square will be much more sensitive to the small-scale clumpiness of the gas maps. Finally, the diffusion results in blue show that our model is able to learn a much stronger correlation across all scales. While we are able to achieve high correlations up to around $k=4$ for dark matter, below this scale we lack sufficient constraining power on the substructure from the observables. To ameliorate this, the diffusion model learns to suggest plausible substructures which, while not exactly the true configuration observed in our simulations, are physically consistent with the larger structures. Thus, sometimes these substructures will be correct, and at other times they may be wrong, and the overall net effect will reduce the cross-correlations at smaller scales. de Andres et al. (2024) reconstructs the generated and simulated power spectrum and shows agreement by comparing the relative error, but does not compute $b(k)$ or $r(k)$, which we focus on to show the improvements of our model over linear models in the spectral domain.
In Table 1, we summarize the average values of Figures 8 and 9. Our bias values are close in agreement with unity across all scales for both gas and dark matter maps. Our cross-correlation values are close to unity at large scales, and around $0.97$ for gas and $0.80$ for dark matter for small scales, indicating that our model is able to reproduce the structure of clusters very well at these two different scale lengths.
@@
high mass Gas $0.974_{-0.003}^{+0.004}$ $0.988_{-0.007}^{+0.009}$ $0.970_{-0.006}^{+0.008}$ $0.994_{-0.005}^{+0.004}$
Dark Matter $0.828_{-0.106}^{+0.096}$ $0.987_{-0.010}^{+0.007}$ $0.791_{-0.137}^{+0.125}$ $0.982_{-0.015}^{+0.015}$
low mass Gas $0.982_{-0.003}^{+0.005}$ $0.989_{-0.005}^{+0.004}$ $0.975_{-0.007}^{+0.008}$ $0.995_{-0.004}^{+0.003}$
Dark Matter $0.860_{-0.087}^{+0.077}$ $0.972_{-0.008}^{+0.009}$ $0.831_{-0.114}^{+0.101}$ $0.986_{-0.011}^{+0.012}$
@@ 5 Conclusion
Our diffusion model is able to generate highly realistic and accurate realizations of gas and dark matter maps. The samples from the model converge since a ten-sample mean is able to achieve over one e-folding of a two-sample variance for our test dataset. In addition, because the dataset is normalized by a global maximum and minimum so as to preserve relative intensities between clusters, the model is able to accurately predict the matter distribution across differently-sized clusters. Another indication that our model is able to distinguish between differently-sized clusters is that it is able to reconstruct the mean and spread of both the gas and dark matter density profiles to within 5% on all spatial scales and all cluster sizes. Finally, in the spectral domain, the model achieves a close-to-unity agreement for the bias coefficients and high cross-correlation coefficients between the simulated and generated density maps, implying that our model learns a strong nonlinear mapping that can accurately probe both large and small scale cluster structures.
@@
Future work can be done to improve the reconstruction of the mass maps in both the spatial and spectral domains. Currently, our loss is only the score-matching term, which does not include any information on the physical constraints of our data. We can utilize physics-informed losses during learning, where we augment our loss with extra terms, such as constraints on the total mass of the cluster or regularization terms on the bias and cross-correlation coefficients at different scale lengths. Improvements can also be made on the data end: by adding realistic and structured noise to our dataset, we can move towards using our score models to predict unknown density maps using observational X-ray and SZ images. The noising procedure of the score models will have to be subsequently adapted such that $\sigma_{min}$ matches the level of the inherent data noise, and consequently we expect our constraining power to be worse when we add noise, allowing us only to focus on high signal-to-noise systems and recovering larger-scale features.
Additional work can be done to analyze the dependency of these diffusion models on different hydrodynamical simulations. Since our mock observations of halos and baryon fractions are cosmology dependent, future work will to be done to test the diffusion model dependence on cosmological parameters. Finally, to shift our models towards real observations, we plan to train on data generated from simulations with AGN feedback such as IllustrisTNG and The Three Hundred Projects (e.g. Nelson et al., 2019; Cui et al., 2018), as feedback may weaken the correlation between the densities and the observed SZ and X-ray measurements. We also plan to use weak lensing measurements as an additional conditional measurement: since it is largely AGN independent as it depends on total matter, most of which is dark matter, it will act as a feedback regularizer in the diffusion model training.
@@
The success of our diffusion model to accurately predict gas and dark matter realizations by sampling from a learned distribution has huge implications in real-data inference: these diffusion models can ultimately be used to predict unknown gas and dark matter distributions of real galaxy clusters of interest.
Acknowledgements.
@@
We thank Jun-Yan Zhu for discussions on generative modeling, and Nick Gnedin, Yuuki Omori, and Qirong Zhu for discussions on HYPER. We are grateful to Alexandre Adam for making his implementation of score models publicly available. We thank Jean-Baptiste Melin for providing cluster pressure profiles, and Erwin Lau and Daisuke Nagai for the tabulated cluster profiles from the TNG300 simulation. A.H. thanks Ashley Villar for guidance on parts of this project during the Spring 2024 ASTRON205 course at Harvard University. The HYPER simulation was run on Endeavour at the NASA Advanced Supercomputing (NAS) Division. The machine learning work was run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University. This project is supported by NASA grant 80NSSC22K0821.
Appendix A Halo Profiles
Figure 10: The average (cyan) pressure profile and $1\sigma$ dispersion (shaded) for a subsample of galaxy clusters from the HYPER simulation with $M_{\mathrm{200m}}\approx 6\times 10^{14}\,M_{\odot}$ and $M_{\mathrm{200c}}\approx 4\times 10^{14}\,M_{\odot}$. The simulated profiles agree with the DPP (orange; He et al., 2021) to within $\sim 20\%$. Additional, there is good agreement with the best-fit profiles for the low-mass (dotted) and high-mass (dashed) subsamples from Melin & Pratt (2023).
The HYPER simulation uses the debiased pressure profile (DPP; He et al., 2021) in the subgrid model for the ICM. To compare the simulation with the model, we use a subsample of halos positioned between our low-mass and high-mass ranges. Note that HYPER outputs radial profiles based on the halo mass $M_{\mathrm{200c}}$ within radius $R_{\mathrm{200c}}$, where the average density is 200 times the critical density. The pressure has been normalized by the characteristic pressure (e.g. Lau et al., 2015),
$P_{\mathrm{200c}}=\frac{GM_{\mathrm{200c}}}{2R_{\mathrm{200c}}}(200\rho_{\mathrm{crit}})f_{\mathrm{b}},$ (18)
where $f_{\mathrm{b}}=\Omega_{\mathrm{b}}/\Omega_{\mathrm{m}}$ is the cosmic baryon fraction. Figure 10 shows that the simulated pressure profiles agree with the DPP to within $\sim 20\%$ over the radial range of interest. For comparison, we also show the best-fit profiles for the low-mass and high-mass subsamples from Melin & Pratt (2023), based on joint measurement from Planck and SPT-SZ.
Figure 11: The average (solid) density profile and $1\sigma$ dispersion (shaded) for a subsample of galaxy clusters from the HYPER simulation with $M_{\mathrm{200m}}\approx 6\times 10^{14}\,M_{\odot}$ and $M_{\mathrm{200c}}\approx 4\times 10^{14}\,M_{\odot}$. The dark matter (cyan) density is in excellent agreement with the NFW profile (dashed). The gas density (orange) shows overall good agreement with the TNG300 simulation (gray) despite the different ICM modeling.
Figure 11 shows the density profiles for the same subsample of simulated galaxy clusters. The dark matter density is in excellent agreement with the NFW universal density profile (Navarro et al., 1997). The gas density is suppressed relative to the dark matter at small radii due to pressure support. Despite the different ICM modeling, there is also overall good agreement with the TNG300 simulation (e.g. Barnes et al., 2018). The HYPER gas densities are slightly lower at small radii, likely due to having higher central pressure.